# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

### v1

@@ B.2. Classification results
C25 Electrolytic or electrophoretic processes; apparatus therefor 1.56 1.38
C30 Crystal growth 2.12 1.53
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 0.66 0.23
C99 Subject matter not otherwise provided for in section c 2.8 0.0
D01 Natural or man-made threads or fibres; spinning 0.44 0.35
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 0.22 0.13
D03 Weaving 0.16 0.11
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 0.35 0.24
D05 Sewing; embroidering; tufting 0.07 0.02
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 0.38 0.27
D07 Ropes; cables other than electric 0.13 0.07
D10 Indexing scheme associated with sublasses of section d, relating to textiles 0.44 0.32
D21 Paper-making; production of cellulose 0.85 1.08
D99 Subject matter not otherwise provided for in section d 0.7 3.31
E01 Construction of roads, railways, or bridges 0.6 0.57
E02 Hydraulic engineering; foundations; soil shifting 1.13 0.83
E03 Water supply; sewerage 3.4 2.69
E04 Building 1.18 2.2
E05 Locks; keys; window or door fittings; safes 0.18 0.16
E06 Doors, windows, shutters, or roller blinds, in general; ladders 0.58 1.2
@@
E21 Earth drilling; mining 0.48 0.3
@@
F01 Machines or engines in general; engine plants in general; steam engines 2.64 1.34
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 3.03 1.15
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 4.84 4.87
F04 Engines or pumps 0.48 0.33
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 3.18 1.12
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 0.29 0.15
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 0.41 0.26
F17 Storing or distributing gases or liquids 0.92 0.5
F21 Lighting 0.98 1.62
F22 Steam generation 2.03 3.17
F23 Combustion apparatus; combustion processes 2.92 3.06
F24 Heating; ranges; ventilating 2.03 3.88
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 1.3 1.59
F26 Drying 0.77 0.91
F27 Furnaces; kilns; ovens; retorts 1.76 2.18
F28 Heat exchange in general 1.53 1.76
F41 Weapons 0.12 0.04
F42 Ammunition; blasting 0.2 0.07
@@
G03 Photography; cinematography; analogous techniques using waves other than optical waves 0.17 0.08
G04 Horology 0.25 0.19
G05 Controlling; regulating 1.05 0.73
G06 Computing; calculating; counting 0.52 0.25
G07 Checking-devices 0.3 0.28
G08 Signalling 0.59 0.5
G09 Educating; cryptography; display; advertising; seals 0.33 0.28
G10 Musical instruments; acoustics 0.28 0.19
G11 Information storage 0.24 0.04
G12 Instrument details 0.22 0.12
G16 Information and communication technology [ict] specially adapted for specific application fields 0.57 0.22
G21 Nuclear physics; nuclear engineering 3.16 0.83
@@
H01 Electric elements 1.78 1.36
H02 Generation, conversion, or distribution of electric power 2.01 3.52
H03 Electronic circuitry 0.35 0.08
@@
H05 Electric techniques not otherwise provided for 1.01 0.64
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 1.88 1.25
H99 Subject matter not otherwise provided for in section H 0.76 1.2
Y02 Technologies or applications for mitigation or adaptation against climate change 6.92 7.52
Y04 Information or communication technologies having an impact on other technology areas 5.62 4.75
Y10 Technical subjects covered by former uspc 0.96 0.8

### v2

@@ B.2. Classification results
C25 Electrolytic or electrophoretic processes; apparatus therefor 1.56 1.44
C30 Crystal growth 2.12 1.11
C40 Combinatorial chemistry; libraries, e.g. chemical libraries, dna libraries 0.65 0.12
C99 Subject matter not otherwise provided for in section c 2.79 0.0
D01 Natural or man-made threads or fibres; spinning 0.44 0.39
D02 Yarns; mechanical finishing of yarns or ropes; warping or beaming 0.21 0.14
D03 Weaving 0.16 0.09
D04 Braiding; lace-making; knitting; trimmings; non-woven fabrics 0.34 0.19
D05 Sewing; embroidering; tufting 0.07 0.01
D06 Treatment of textiles or the like; laundering; flexible materials not otherwise provided for 0.38 0.3
D07 Ropes; cables other than electric 0.12 0.05
D10 Indexing scheme associated with sublasses of section d, relating to textiles 0.44 0.31
D21 Paper-making; production of cellulose 0.85 1.15
D99 Subject matter not otherwise provided for in section d 0.62 3.21
E01 Construction of roads, railways, or bridges 0.59 0.61
E02 Hydraulic engineering; foundations; soil shifting 1.13 0.86
E03 Water supply; sewerage 3.39 2.72
E04 Building 1.18 1.07
E05 Locks; keys; window or door fittings; safes 0.18 0.12
E06 Doors, windows, shutters, or roller blinds, in general; ladders 0.58 0.57
@@
E21 Earth drilling; mining 0.48 0.23
@@
F01 Machines or engines in general; engine plants in general; steam engines 2.63 1.57
F02 Internal-combustion engines; hot-gas or combustion-product engine plants 3.02 1.35
F03 Machines or engines for liquids; wind, spring, weight, or miscellaneous motors 4.82 5.36
F04 Engines or pumps 0.48 0.37
F05 Indexing schemes relating to engines or pumps in various subclasses of classes f01-f04 3.17 1.26
F15 Fluid-pressure actuators; hydraulics or pneumatics in general 0.29 0.18
F16 Engineering elements or units; general measures for effecting and maintaining effective functioning of machines 0.41 0.27
F17 Storing or distributing gases or liquids 0.92 0.37
F21 Lighting 0.98 1.76
F22 Steam generation 2.02 3.64
F23 Combustion apparatus; combustion processes 2.91 3.77
F24 Heating; ranges; ventilating 2.03 4.12
F25 Refrigeration or cooling; combined heating and refrigeration systems; heat pump systems 1.3 1.69
F26 Drying 0.77 1.04
F27 Furnaces; kilns; ovens; retorts 1.75 2.57
F28 Heat exchange in general 1.53 1.88
F41 Weapons 0.11 0.04
F42 Ammunition; blasting 0.2 0.06
@@
G03 Photography; cinematography; analogous techniques using waves other than optical waves 0.17 0.06
G04 Horology 0.25 0.21
G05 Controlling; regulating 1.05 0.72
G06 Computing; calculating; counting 0.51 0.2
G07 Checking-devices 0.3 0.29
G08 Signalling 0.59 0.55
G09 Educating; cryptography; display; advertising; seals 0.33 0.27
G10 Musical instruments; acoustics 0.28 0.12
G11 Information storage 0.24 0.03
G12 Instrument details 0.22 0.11
G16 Information and communication technology [ict] specially adapted for specific application fields 0.56 0.15
G21 Nuclear physics; nuclear engineering 3.15 0.54
@@
H01 Electric elements 1.77 1.26
H02 Generation, conversion, or distribution of electric power 2.01 3.92
H03 Electronic circuitry 0.35 0.07
@@
H05 Electric techniques not otherwise provided for 1.01 0.61
H10 Semiconductor devices; electric solid-state devices not otherwise provided for 2.25 2.66
H99 Subject matter not otherwise provided for in section H 0.7 1.2
Y02 Technologies or applications for mitigation or adaptation against climate change 6.88 7.55
Y04 Information or communication technologies having an impact on other technology areas 5.6 4.84
Y10 Technical subjects covered by former uspc 0.96 0.84